# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

### v4

@@
Abstract
Integrating powerful but computationally expensive Pre-trained Language Models (PLMs) with Graph Neural Networks (GNNs) is a key challenge, especially on text-rich heterophilic graphs. We propose the Graph Masked Language Model (GMLM), a framework designed for the efficient and effective fusion of graph structure and text semantics. GMLM employs a two-stage process: first, a contrastive pre-training stage with a novel soft masking technique builds a robust multi-scale GNN; second, an end-to-end fine-tuning stage uses a dynamic active node selection strategy for scalability and a bi-directional cross-attention module for deep fusion. Experiments on five heterophilic benchmarks show GMLM achieves state-of-the-art results on four, significantly outperforming prior GNN and large LLM-based methods. For instance, it improves accuracy on the Texas dataset by over 8% and on Wisconsin by nearly 5%. Our work demonstrates that a sophisticated, deeply-integrated architecture can be more effective and efficient than larger, general-purpose models for text-rich graph representation learning.
GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification
Aarush Sinha Independent Researcher aarush.sinha@gmail.com
1 Introduction
@@
Graph Neural Networks (GNNs) Scarselli et al. (2009); Zhou et al. (2021) have demonstrated remarkable success in node classification, link prediction, and graph-level tasks. However, many such architectures focus primarily on structural information and local neighborhood features, potentially overlooking rich semantic content embedded within node attributes, especially textual attributes. While recent advances in natural language processing, particularly Pre-trained Language Models (PLMs) like BERT Devlin et al. (2019), have revolutionized our ability to capture contextual information from text, their effective and scalable integration with graph learning remains an active area of research. The challenge is particularly pronounced in heterophilic graphs, where connected nodes may exhibit dissimilar features and labels, requiring a nuanced understanding of both structure and semantics.
@@
While architectures like Graph Attention Networks (GATs) Veličković et al. (2018) offer learnable attention mechanisms for neighborhood aggregation, and Relational Graph Convolutional Networks (RGCNs) Schlichtkrull et al. (2017) can handle typed edges, they do not inherently perform deep semantic reasoning on node-associated texts without explicit integration with PLMs. Concurrently, masked modeling techniques, central to the success of PLMs, have proven highly effective for learning contextual representations. This prompts the question: How can we develop a deeply integrated framework that combines the strengths of GNNs and PLM-style masked modeling to enhance representation learning on text-rich graphs, while addressing computational scalability?
@@
Traditional node feature masking in GNNs Mishra et al. (2021) often involves random feature perturbation or complete removal, which can disrupt learning or fail to provide sufficiently informative training signals. Furthermore, applying PLMs to all node texts in large graphs is often computationally prohibitive.
@@
In this paper, we present a novel framework that deeply integrates GNNs and PLMs for improved node representation learning on text-rich graphs, with a focus on heterophilic settings and computational efficiency. Our approach introduces three key innovations:
@@
This combination of selective PLM processing, robust GNN pre-training, and sophisticated fusion allows our model to learn both structural and semantic patterns effectively, leading to more powerful and contextually aware node representations for challenging node classification tasks.
Our main contributions in this work can be summarized as follows:
  1. 1.
    We propose a novel and scalable framework that deeply integrates a multi-scale GNN with a PLM. The framework’s efficiency is achieved through a dynamic active node selection mechanism for on-demand text processing, making it tractable for large graphs.
  2. 2.
    We introduce a GNN-specific contrastive pre-training stage that utilizes a soft feature masking technique with a learnable graph [MASK] token. This stage is designed to effectively train a deep, multi-scale GNN architecture, allowing it to learn robust and hierarchical structural representations before end-to-end fine-tuning (Section 5.1).
  3. 3.
    We design a sophisticated bi-directional cross-attention fusion module that moves beyond simple concatenation. This mechanism allows the graph and text representations to mutually inform and refine one another, creating a richer, more contextually-aware joint representation for improved node classification, especially on challenging heterophilic graphs (Section 5.2).
2 Related Works
Graph Neural Networks (GNNs) Zhou et al. (2021) provide a powerful framework for learning from graph-structured data. Building upon early spectral and spatial convolutional methods, Graph Attention Networks (GATs) Veličković et al. (2018) integrated self-attention Vaswani et al. (2017). This allows adaptive weighting of neighbors during message passing, enhancing model expressivity and suitability for irregular graph structures, as detailed in recent surveys.
Efforts to adapt the Transformer architecture, initially designed for sequences, to graph data aim to capture long-range dependencies via global self-attention, potentially overcoming limitations of traditional message passing Ying et al. (2021), Yun et al. (2020). These Graph Transformers, reviewed systematically Shehzad et al. (2024), can incorporate positional information, handle diverse graph scales, achieve competitive results, and utilize language model pretraining benefits on text-rich graphs.
Masked Pretraining Strategies, popularized by language models Devlin et al. (2019) that learn contextual representations via masked token prediction, have been adapted for graphs. The Masked Graph Autoencoder (MaskGAE) Li et al. (2023a) framework, for instance, masks graph edges and trains the model to reconstruct them, thereby learning robust representations for downstream tasks like link prediction and node classification.
Integration of Language Models and Graph Structures seeks to combine textual and structural information. Key directions include: fusing pretrained language models (LMs) with GNNs for joint reasoning over text and structure Plenz and Frank (2024); injecting structural information directly into LMs using graph-guided attention mechanisms, potentially removing the need for separate GNN modules Yuan and Färber (2024); interpreting Transformers as GNNs operating on complete graphs, allowing explicit modeling of edge information Henderson et al. (2023); and jointly training large LMs and GNNs to utilize both contextual understanding and structural learning capabilities Ioannidis et al. (2022).
3 Preliminaries
@@
3.1 Graph Representation and Notation
A graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of $N=|\mathcal{V}|$ nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$ is the set of $M=|\mathcal{E}|$ edges. Each edge $(v_{i},r,v_{j})\in\mathcal{E}$ connects node $v_{i}$ to node $v_{j}$ under a relation type $r\in\mathcal{R}$, where $\mathcal{R}$ is the set of possible relation types. For graphs without explicit multiple relation types, we can consider a single relation type, effectively making it $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Each node $v_{i}\in\mathcal{V}$ is associated with an initial feature vector $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$, where $d_{x}$ is the input feature dimensionality. In the context of this work, nodes are also associated with textual descriptions $T=\{t_{1},t_{2},\dots,t_{N}\}$, where $t_{i}$ is the raw text for node $v_{i}$. The goal of node classification is to predict a label $y_{i}\in\mathcal{Y}$ for each node (or a subset of nodes), where $\mathcal{Y}$ is the set of possible labels.
3.2 Graph Neural Networks (GNNs)
Graph Neural Networks (GNNs) are a class of neural networks designed to operate directly on graph-structured data. They typically learn node representations by aggregating information from their local neighborhoods. A generic GNN layer can be expressed as:
$\mathbf{h}_{i}^{(l+1)}=\text{UPDATE}^{(l)}\left(\mathbf{h}_{i}^{(l)},\text{AGGREGATE}^{(l)}\left(\{(\mathbf{h}_{j}^{(l)},\mathbf{e}_{ij}):v_{j}\in\mathcal{N}(v_{i})\}\right)\right)$
where $\mathbf{h}_{i}^{(l)}$ is the representation of node $v_{i}$ at layer $l$, $\mathcal{N}(v_{i})$ is the set of neighbors of $v_{i}$, $\mathbf{e}_{ij}$ is the feature of the edge between $v_{i}$ and $v_{j}$ (if available), and AGGREGATE and UPDATE are differentiable functions ( sum, mean, max, or more complex neural networks).
3.2.1 Relational Graph Convolutional Networks (RGCNs)
Relational Graph Convolutional Networks (RGCNs) Schlichtkrull et al. (2017) extend GCNs to handle multi-relational graphs (heterogeneous graphs with different edge types). The message passing for a node $v_{i}$ in an RGCN layer is typically defined as:
$\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{v_{j}\in\mathcal{N}_{r}(v_{i})}\frac{1}{c_{i,r}}\mathbf{W}_{r}^{(l)}\mathbf{h}_{j}^{(l)}+\mathbf{W}_{0}^{(l)}\mathbf{h}_{i}^{(l)}\right)$ (2)
@@ 3.2.1 Relational Graph Convolutional Networks (RGCNs)
3.3 Pre-trained Language Models (PLMs)
@@
Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2019) and its variants ( DistilBERT Sanh et al. (2020)), have achieved state-of-the-art results on various natural language processing (NLP) tasks. These models are typically based on the Transformer architecture Vaswani et al. (2023) and are pre-trained on massive text corpora using self-supervised objectives like Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). Given an input text sequence ( a node’s textual description $t_{i}$), a PLM first tokenizes it into a sequence of sub-word tokens $[w_{1},w_{2},\dots,w_{S}]$. These tokens are then converted into input embeddings (sum of token, position, and segment embeddings) and fed into multiple Transformer layers. The output of the PLM is a sequence of contextualized embeddings for each token. For sentence-level or document-level tasks, often the embedding of a special [CLS] token ( $\mathbf{h}_{[CLS]}$) or an aggregation ( mean pooling) of token embeddings is used as the representation of the entire input text. In our work, we utilize a PLM (specifically DistilBERT) to obtain semantic embeddings $\mathbf{h}_{PLM}(t_{i})$ from node texts $t_{i}$.
3.4 Node Feature Masking in GNNs
@@
Masking is a common technique in self-supervised learning. In the context of GNNs, node feature masking aims to train the model to reconstruct or predict properties of masked nodes or their features, thereby learning robust representations. A simple approach might involve replacing a fraction of node features $\mathbf{x}_{i}$ with a zero vector or a special [MASK] token embedding before feeding them into the GNN Hu et al. (2020); You et al. (2020). Our GNN contrastive pretraining phase builds upon this idea but introduces a soft masking mechanism with a learnable graph-specific [MASK] token.
3.5 Heterophily in Graphs
@@
Graph homophily refers to the principle that nodes tend to connect to other nodes that are similar to themselves ( in terms of features or labels). Conversely, graph heterophily (or disassortativity) describes the scenario where connected nodes are often dissimilar Zhu et al. (2020a). Many real-world graphs exhibit heterophily, such as protein-protein interaction networks or certain social networks. Standard GNNs, which assume homophily by smoothing features over neighborhoods, may perform poorly on heterophilic graphs. Designing GNN architectures that can effectively handle heterophily is an ongoing research challenge, often requiring models to capture more complex relational patterns beyond simple neighborhood similarity. Our work aims to contribute to this by integrating rich textual semantics which can provide crucial context in such settings.
3.6 Contrastive Learning
@@
Contrastive learning is a self-supervised learning paradigm that aims to learn representations by maximizing agreement between differently augmented "views" of the same data sample (positive pairs) while minimizing agreement between views of different samples (negative pairs). A common contrastive loss function is the NT-Xent loss Sohn (2016):
$\mathcal{L}_{NT\text{-}Xent}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1}^{2B}\mathbb{I}_{[k\neq i]}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
@@
Figure 1: Caption
4 Methodology
We introduce a novel architecture designed to effectively integrate deep graph structural information with rich textual semantics from nodes for robust node representation learning. The architecture is characterized by two main processing pathways for graph and text data, a multi-scale GNN with residual connections, a sophisticated bi-directional cross-attention fusion module, and specialized input perturbation mechanisms to enhance robustness and training stability.
4.1 GNN Input Feature Perturbation with Soft Masking
@@
$\mathbf{x}^{\prime}_{i}=\begin{cases}(1-\beta)\cdot\mathbf{x}_{i}+\beta\cdot\mathbf{e}_{mask}^{gnn}&\text{if }m_{pert,i}=1\\ \mathbf{x}_{i}&\text{if }m_{pert,i}=0\end{cases}$ (4)
@@
4.2 Multi-Scale Relational GNN Branch
@@
$\mathbf{Z}^{(l)}$ $=\text{GELU}(\text{GraphNorm}(\text{RGCN}^{(l)}(\mathbf{H}^{(l-1)},\mathcal{E})))$ (5)
$\mathbf{H}^{(l)}$ $=\text{Dropout}(\mathbf{Z}^{(l)})+\text{Proj}^{(l-1)}(\mathbf{H}^{(l-2)})$ (6)
@@
$\mathbf{h}_{G,i}=\text{LayerNorm}\left(\sum_{l=1}^{4}w_{l}\cdot\text{Proj}^{(l)}(\mathbf{h}_{i}^{(l)})\right)$ (7)
@@
4.3 Pre-trained Language Model (PLM) Branch
@@
For each node $v_{i}\in\mathcal{V}_{active}$, its text $t_{i}$ is tokenized and fed into a pre-trained language model, such as GTE-Base (110M)Li et al. (2023b) and Snowflake-Embed(305M)Yu et al. (2024). To derive a fixed-size representation from the PLM’s variable-length output, we perform mean pooling over the last hidden state, weighted by the attention mask to exclude padding tokens. The PLM embedding for an active node $v_{i}$ is:
$\mathbf{h}_{PLM,i}=\frac{\sum_{j=1}^{L}\mathbf{o}_{j}\cdot a_{j}}{\sum_{j=1}^{L}a_{j}}$ (8)
@@
4.4 Bi-Directional Attention Fusion and Classification

### v6

@@
Abstract.
Integrating Pre-trained Language Models (PLMs) with Graph Neural Networks (GNNs) remains a central challenge in text-rich heterophilic graph learning. We propose a novel integration framework that enables effective fusion between powerful pre-trained text encoders and Relational Graph Convolutional networks (R-GCNs). Our method enhances the alignment of textual and structural representations through a bi-directional fusion mechanism and contrastive node-level optimization. To evaluate the approach, we train two variants using different PLMs: Snowflake-Embed (state-of-the-art) and GTE-base each paired with an R-GCN backbone. Experiments on five heterophilic benchmarks demonstrate that our integration method achieves state-of-the-art results on four datasets, surpassing existing GNN and large language model–based approaches. Notably, Snowflake-Embed + R-GCN improves accuracy on the Texas dataset by over 8% and on Wisconsin by nearly 5%. These results highlight the effectiveness of our fusion strategy for advancing text-rich graph representation learning.
Heterophilic Node Classification, RGCN, Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning
1. Introduction
@@
Graph Neural Networks (GNNs) (Scarselli et al., 2009; Zhou et al., 2021) have demonstrated remarkable success in node classification, link prediction, and graph-level tasks. However, many such architectures focus primarily on structural information and local neighborhood features, potentially overlooking rich semantic content embedded within node attributes, especially textual attributes. While recent advances in natural language processing, particularly Pre-trained Language Models (PLMs) like BERT (Devlin et al., 2019), have revolutionized our ability to capture contextual information from text, their effective and scalable integration with graph learning remains an active area of research. The challenge is particularly pronounced in heterophilic graphs, where connected nodes may exhibit dissimilar features and labels, requiring a nuanced understanding of both structure and semantics.
@@
While architectures like Graph Attention Networks (GATs) (Veličković et al., 2018) offer learnable attention mechanisms for neighborhood aggregation, and Relational Graph Convolutional Networks (RGCNs) (Schlichtkrull et al., 2017) can handle typed edges, they do not inherently perform deep semantic reasoning on node-associated texts without explicit integration with PLMs. Concurrently, masked modeling techniques, central to the success of PLMs, have proven highly effective for learning contextual representations. This prompts the question: How can we develop a deeply integrated framework that combines the strengths of GNNs and PLM-style masked modeling to enhance representation learning on text-rich graphs ?
@@
Traditional node feature masking in GNNs (Mishra et al., 2021) often involves random feature perturbation or complete removal, which can disrupt learning or fail to provide sufficiently informative training signals. Furthermore, applying PLMs to all node texts in large graphs is often computationally prohibitive.
@@
In this paper, we present a novel framework that deeply integrates GNNs and PLMs for improved node representation learning on text-rich graphs, with a focus on heterophilic settings. Our approach introduces three key innovations:
@@
2. Related Works
Graph Neural Networks (GNNs) (Zhou et al., 2021) provide a powerful framework for learning from graph-structured data. Building upon early spectral and spatial convolutional methods, Graph Attention Networks (GATs) (Veličković et al., 2018) integrated self-attention (Vaswani et al., 2017). This allows adaptive weighting of neighbors during message passing, enhancing model expressivity and suitability for irregular graph structures, as detailed in recent surveys.
Efforts to adapt the Transformer architecture, initially designed for sequences, to graph data aim to capture long-range dependencies via global self-attention, potentially overcoming limitations of traditional message passing (Ying et al., 2021), (Yun et al., 2020). These Graph Transformers, reviewed systematically (Shehzad et al., 2024), can incorporate positional information, handle diverse graph scales, achieve competitive results, and utilize language model pretraining benefits on text-rich graphs.
Masked Pretraining Strategies, popularized by language models (Devlin et al., 2019) that learn contextual representations via masked token prediction, have been adapted for graphs. The Masked Graph Autoencoder (MaskGAE) (Li et al., 2023a) framework, for instance, masks graph edges and trains the model to reconstruct them, thereby learning robust representations for downstream tasks like link prediction and node classification.
Integration of Language Models and Graph Structures seeks to combine textual and structural information. Key directions include: fusing pretrained language models (LMs) with GNNs for joint reasoning over text and structure (Plenz and Frank, 2024); injecting structural information directly into LMs using graph-guided attention mechanisms, potentially removing the need for separate GNN modules (Yuan and Färber, 2024); interpreting Transformers as GNNs operating on complete graphs, allowing explicit modeling of edge information (Henderson et al., 2023); and jointly training large LMs and GNNs to utilize both contextual understanding and structural learning capabilities (Ioannidis et al., 2022).
Figure 1. The text related to each node is fed into the text-encoder and the node features by the RGCN layers that use multi-layer embeddings fused together. The Cross Attention layers between Graph and Text are then concatenated by a fusion layer which is finally trained and evaluated.
3. Preliminaries
@@
3.1. Relational Graph Convolutional Networks (RGCNs)
Relational Graph Convolutional Networks (RGCNs) (Schlichtkrull et al., 2017) extend GCNs to handle multi-relational graphs (heterogeneous graphs with different edge types). The message passing for a node $v_{i}$ in an RGCN layer is typically defined as:
(1) $\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{v_{j}\in\mathcal{N}_{r}(v_{i})}\frac{1}{c_{i,r}}\mathbf{W}_{r}^{(l)}\mathbf{h}_{j}^{(l)}+\mathbf{W}_{0}^{(l)}\mathbf{h}_{i}^{(l)}\right)$
@@ 3.1. Relational Graph Convolutional Networks (RGCNs)
3.2. Pre-trained Language Models (PLMs)
@@
Pre-trained Language Models (PLMs), such as BERT (Devlin et al., 2019) and its variants (DistilBERT (Sanh et al., 2020)), have achieved state-of-the-art results on various natural language processing (NLP) tasks. These models are typically based on the Transformer architecture (Vaswani et al., 2023) and are pre-trained on massive text corpora using self-supervised objectives like Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). Given an input text sequence ( a node’s textual description $t_{i}$), a PLM first tokenizes it into a sequence of sub-word tokens $w_{1},w_{2},\dots,w_{S}]$. These tokens are then converted into input embeddings (sum of token, position, and segment embeddings) and fed into multiple Transformer layers. The output of the PLM is a sequence of contextualized embeddings for each token. For sentence-level or document-level tasks, often the embedding of a special [CLS] token ( $\mathbf{h}_{[CLS]}$) or an aggregation ( mean pooling) of token embeddings is used as the representation of the entire input text. In our work, we utilize a PLM to obtain semantic embeddings $\mathbf{h}_{PLM}(t_{i})$ from node texts $t_{i}$.
3.3. Node Feature Masking in GNNs
@@
Masking is a common technique in self-supervised learning. In the context of GNNs, node feature masking aims to train the model to reconstruct or predict properties of masked nodes or their features, thereby learning robust representations. A simple approach might involve replacing a fraction of node features $\mathbf{x}_{i}$ with a zero vector or a special [MASK] token embedding before feeding them into the GNN (Hu et al., 2020; You et al., 2020). Our GNN contrastive pretraining phase builds upon this idea but introduces a soft masking mechanism with a learnable graph-specific [MASK] token.
3.4. Heterophily in Graphs
@@
Graph homophily refers to the principle that nodes tend to connect to other nodes that are similar to themselves ( in terms of features or labels). Conversely, graph heterophily (or disassortativity) describes the scenario where connected nodes are often dissimilar (Zhu et al., 2020a). Many real-world graphs exhibit heterophily, such as protein-protein interaction networks or certain social networks. Standard GNNs, which assume homophily by smoothing features over neighborhoods, may perform poorly on heterophilic graphs. Designing GNN architectures that can effectively handle heterophily is an ongoing research challenge, often requiring models to capture more complex relational patterns beyond simple neighborhood similarity. Our work aims to contribute to this by integrating rich textual semantics which can provide crucial context in such settings.
3.5. Contrastive Learning
@@
Contrastive learning is a self-supervised learning paradigm that aims to learn representations by maximizing agreement between differently augmented "views" of the same data sample (positive pairs) while minimizing agreement between views of different samples (negative pairs). A common contrastive loss function is the NT-Xent loss (Sohn, 2016):
(2) $\mathcal{L}_{NT\text{-}Xent}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1}^{2B}\mathbb{I}_{[k\neq i]}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
@@
4. Methodology
We introduce a novel architecture designed to effectively integrate deep graph structural information with rich textual semantics from nodes for robust node representation learning. The architecture is characterized by two main processing pathways for graph and text data, a multi-scale GNN with residual connections, a sophisticated bi-directional cross-attention fusion module, and specialized input perturbation mechanisms to enhance robustness and training stability. We provide an overview of our proposed architecture in Figure 1. All experiments were done on a single RTX8000(45GB) using PyTorch (Paszke et al., 2019) and PyG(Fey et al., 2025)(Fey and Lenssen, 2019).
4.1. GNN Input Feature Perturbation with Soft Masking
@@
(3) $\mathbf{x}^{\prime}_{i}=\begin{cases}(1-\beta)\cdot\mathbf{x}_{i}+\beta\cdot\mathbf{e}_{mask}^{gnn}&\text{if }m_{pert,i}=1\\ \mathbf{x}_{i}&\text{if }m_{pert,i}=0\end{cases}$
@@
4.2. Multi-Scale Relational GNN Branch
@@
(4) $\mathbf{Z}^{(l)}$ $=\text{GELU}(\text{GraphNorm}(\text{RGCN}^{(l)}(\mathbf{H}^{(l-1)},\mathcal{E})))$
(5) $\mathbf{H}^{(l)}$ $=\text{Dropout}(\mathbf{Z}^{(l)})+\text{Proj}^{(l-1)}(\mathbf{H}^{(l-2)})$
@@
(6) $\mathbf{h}_{G,i}=\text{LayerNorm}\left(\sum_{l=1}^{4}w_{l}\cdot\text{Proj}^{(l)}(\mathbf{h}_{i}^{(l)})\right)$
@@
4.3. Pre-trained Language Model (PLM) Branch
@@
For each node $v_{i}\in\mathcal{V}_{active}$, its text $t_{i}$ is tokenized and fed into a pre-trained language model, such as GTE-Base (110M)(Li et al., 2023b) and Snowflake-Embed(305M)(Yu et al., 2024). To derive a fixed-size representation from the PLM’s variable-length output, we perform mean pooling over the last hidden state, weighted by the attention mask to exclude padding tokens. The PLM embedding for an active node $v_{i}$ is:
(7) $\mathbf{h}_{PLM,i}=\frac{\sum_{j=1}^{L}\mathbf{o}_{j}\cdot a_{j}}{\sum_{j=1}^{L}a_{j}}$
@@
4.4. Bi-Directional Attention Fusion and Classification